OpenStudy (anonymous):

(x-7)^2=96 square root property? help please!!!!

OpenStudy (anonymous):

\[(x-7)^2=96\\ x-7=\pm\sqrt{96}\\ x=7\pm\sqrt{96}\]

OpenStudy (anonymous):

you might want to write \(\sqrt{96}\) in simplest radical form